This is a digit problem. I apologize if I am in the wrong area. The question is;

The sum of the digits of the two-digit numeral is 8. The number with the digits interchanged is 7 times the tens digit of the original number. Find the original number
Thanks for your help!
C. Johnson

Answer by stanbon(48558)

The sum of the digits of the two-digit numeral is 8. The number with the digits interchanged is 7 times the tens digit of the original number. Find the original number
-------------
Let the two digit number be tu where t is the "ten's" digit and u and the "unit's digit. Also keep in mind that tu =10t+u just like 23 = 10*2+3.
---------
EQUATIONS:
t+u=8
10u+t = 7t
-------
Solve the 1st for t: t=8-u
Siplify the 2nd then,
Substitute t=8-u into the 2nd:
10u=6t
10u=6(8-u)
10u = 48-6u
16u=48
u=3
The t=8-u=8-3=5
-----------------
The original number is 53
-----------
Cheers,
Stan H.